OpenStudy (anonymous):

Solve for x: log5^x-log5^2=2

OpenStudy (anonymous):

is this \[log_5(x)-log_5(2)=2\]?

OpenStudy (anonymous):

log 5x-log5 with the 2 at the top =2

OpenStudy (anonymous):

if so combine the first two into one log as \[log_5(\frac{x}{2})=2\]\]

OpenStudy (anonymous):

ooh\[log(5) - log(2^2)=2\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

the log5 and the x is superscript

OpenStudy (anonymous):

x = [2 + 2log5]/log5

OpenStudy (anonymous):

did you see it

OpenStudy (anonymous):

yeah...log_5(x/2) = 2, x/2 = 5^2, x = 50

OpenStudy (anonymous):

thank you so much i have to learn how to write it like that

OpenStudy (anonymous):

ah it was the first one i wrote. jamesm has the anwer

OpenStudy (anonymous):

Thank you jamesm